# Observation of a non-Hermitian supersonic mode

### v1

@@ II Crossing the exceptional point
Figure 2: (a) While any many-body wavefunction can be cast as an MPS, it permits especially efficient ground state representations for $1d$ systems. In the right canonical form, each tensor in the MPS is an isometry [104, 135, 136, 137, 108]. To implement on a quantum computer, each isometry is embedded as a $2\chi\times 2\chi$ unitary through a QR decomposition. The MPS is then cast into a sequential quantum circuit. (b) Each $U_{A}$ can be synthesized from SU(4) gates with certain local geometries. Here we demonstrate a ladder geometry and a brickwall geometry. The depth of the local circuit is denoted $\tau$.
@@
Here $E$ is a complex variable we optimize over. When $\mathcal{L}^{\prime}=0$ we are guaranteed that |$\ket{\psi({\bm{\theta}})}$ is a right eigenstate  [^2] For the rest of the paper, we focus on the property of right eigenstates, as the left eigenstates can be found by substituting $H\leftarrow H^{\dagger}$ of $H$ and $E$ is an eigenvalue. Using this method, the authors devised a numerical optimizer and were able to numerically compute the left and right eigenstates, verifying the biorthogonal relations, as well as evaluating many observables [138].
@@
Although the variance minimization scheme presented in Ref. 138 includes all the fundamental components of a variational quantum algorithm, the simulated system sizes were limited to only 7 qubits due to two main obstacles: First, they used a direct, full-state preparation method that is demanding on both classical and quantum memory. Capturing long-range correlations would then require a large circuit volume, making it impractical for larger systems. Second, the termination condition for variance minimization only ensures that some eigenstate is found, but not necessarily the one of interest. Given that the number of eigenvalues increases exponentially with system size, a random guess for $E$ is likely to result in an undesired eigenstate.
@@
Figure 3: (a) The real and (b) imaginary parts of variational right ground state preparation results of a $n=19$ non-Hermitian Ising chain are compared with that from a $\chi=100$ NH-DMRG. The exceptional point can be read from the imaginary part of $E$. We experimentally execute the $J=-1,\kappa\in\{0.2,0.5\}$ circuits for 2,000 measurement shots and report the energy outcomes, which are both within 2% error of the DMRG values. (c) The absolute value of relative energy difference. Notice that the peak on each curve corresponds to the EP read from (b). (d) and (e) repeat the study of (a) and (b), except now the system size is set to 20 so that the system is EP-free. (f) Comparing $ZZ$ correlation functions measured in Quantinuum’s H1 processor (2,000 shots) to DMRG results. All results are prepared with a $q=2$, $\tau=2$ ladder architecture as defined in Fig. 2, and error bars represent one statistical standard deviation.
@@
To solve the first obstacle, we make use of quantum circuit tensor network state (qTNS) techniques [104, 90, 140, 101, 108, 118, 117, 92], to sequentially simulate many-body quantum states, as shown in Fig. 2. Rather than a direct one-to-one encoding between a spin and a qubit, sequential simulation introduces $q$ ancilla qubits, or bond qubits, with local circuit depth $\tau$ to faithfully capture the near-area law entanglement structure of a physical state, generating a sequence of quantum operations that allows one to sample properties of the many-body state along a spatial direction without storing the full state in quantum memory; we defer a more detailed explanation to Appx. A.
@@
For the second obstacle, we designed a ‘warm start’ method, as widely used in many other variational algorithms [116, 118], namely, we first turn off the non-Hermitian field and solve for the Hamiltonian’s ground state with a VQE algorithm and gradually turn on the imaginary field, feeding the previous optimization result as an initialization. Concretely, we consider the $1d$-transversal field Ising model with an imaginary longitudinal field [141]:
@@
where $J$ and $\kappa$ are real numbers. To find the ground state at a certain $\kappa_{\rm targ}$, we execute the procedure in Algo. 1:
Algorithm 2 VQE for dissipative Ising model with variance minimizaiton
1:Set the target Hamiltonian to be $H=H_{\rm Ising}(J,\kappa=0)$
2:Initialize |$\ket{\psi({\bm{\theta}})}$ where the circuit with a variational circuit generated by quantum matrix product state (qMPS) [104, 90, 108]
3:Find the Hamiltonian’s ground state by minimizing $\mathcal{L}=\innerproduct{\psi({\bm{\theta}})|H}{\psi({\bm{\theta}})}$ and record the optimized circuit parameter $\bm{\theta}^{*}$ and energy $E^{*}$;
4:for $\kappa=\Delta\kappa,2\Delta\kappa,\ldots\kappa_{\rm targ}$ do
5:     Initialize the circuit at $\bm{\theta}_{0}\leftarrow\bm{\theta}^{*}$ and $E_{0}\leftarrow E^{*}$
6:     Minimize the target function $\bra{\psi({\bm{\theta}})}(H-E)(H^{\dagger}-E^{*})\ket{\psi({\bm{\theta}})}$ over $\bm{\theta}$ and $E$.
7:     Record optimized parameter $\bm{\theta}^{*}(\kappa)$ and energy $E^{*}(\kappa)$
8:end for
The dissipative Ising model exerts $\mathcal{PT}$-symmetry, but it goes through spontaneous symmetry breaking at $\kappa_{c}$, the so-called exceptional point: at $\kappa<\kappa_{c}$, the ground state energy remains real despite the non-Hermicity; at $\kappa>\kappa_{c}$, the ground state energy merges with the first excited state to form a complex conjugate pair in their energy values. The exceptional point is very sensitive to $J$ and the parity of the length of the chain.
@@
Additional numerical results for a non-Hermitian XXZ chain with size $L=64$ can be found in Appx. B.
@@ III Discussion
We have combined two quantum state compression techniques, GMPS and VQC, to experimentally study the quenched dynamics of an extended strongly-correlated non-Hermitian system. Using a trapped-ion implementation on the Quantinuum H1 quantum processor, we have observed a supersonic mode in the connected density-density correlation function of a fermionic chain following a non-Hermitian quantum quench—a phenomenon conventionally forbidden by the Lieb-Robinson bound in Hermitian systems. The system sizes, time scales, and low sample complexity achieved in our experiments are enabled by the convergence of the advantages offered by universal quantum computers and the variational formulation of the dynamics employed in our experiments, allowing us to bypass costly postselection-based implementations with an affordable approach based on unitary dynamics. Similarly, we exploited the efficient compression offered by qMPS to experimentally prepare eigenstates of a non-hermitian spin chain where we experimentally measured correlation functions and energies across an exceptional point on a dissipative spin chain up to length $n=20$ using only 3 qubits.
@@
Our work raises a number of intriguing questions regarding the opportunities and challenges of using digital quantum computers for simulating non-Hermitian systems. In particular, in Thm. I.1 we established the hardness of the simulation of quench dynamics of certain initial states under single qubit non-Hermitian dynamics. A key open question arising from our experiments and theoretical findings is to determine the properties of the initial state and the non-Hermitian Hamiltonian that enable efficient unitary simulation. Here, a natural direction is to explore the use of VQC to alleviate the postselection requirements which are often encountered in systems such as those exhibiting measurement-induced phase transitions [143, 144, 145]. Similarly, these and other related questions can be reframed in the context of quantum algorithms for solving dissipative differential equations. Specifically, under what conditions can quantum computers provide a significant advantage in solving these equations? This problem has been studied from an algorithmic perspective [146, 147] and can potentially offer new insights into the question of simulability of non-hermitian systems more broadly.
@@
A natural extension of our work is to explore dynamics and eigenstate properties beyond one-dimensional systems where interesting non-Hermitian phenomena emerge [148, 149]. Specifically, asymptotic quantum memory costs to approximate the states in different non-Hermitian phases of matter, as has been done in various Hermitian systems, where one could relate the hardness of compression to entanglement entropy laws [150]. It is worth pursuing what metric one should use in the non-Hermitian case, as there does not exist a good entanglement measure due to the lack of an orthonormal basis [16]. As we push the boundaries of quantum computation, the techniques and insights gained from our study not only illuminate new pathways for exploring non-Hermitian physics but also pave the way for future advancements in quantum simulation aided by the convergence of innovative techniques such as compression and variational compilation with cutting-edge quantum hardware.
IV Acknowledgement

### v2

@@ II.2 Crossing the exceptional point
Figure 2: (a) While any many-body wavefunction can be cast as an MPS, it permits especially efficient ground state representations for $1d$ systems. In the right canonical form, each tensor in the MPS is an isometry [28, 44]. To implement on a quantum computer, each isometry is embedded as a $2\chi\times 2\chi$ unitary through a QR decomposition. The MPS is then cast into a sequential quantum circuit. (b) Each $U_{A}$ can be synthesized from SU(4) gates with certain local geometries. Here we demonstrate a ladder geometry and a brickwall geometry. The depth of the local circuit is denoted $\tau$. The sequential quantum circuit can then be used as a variational ansatz. Due to having limited access to quantum resources, we first classically variationally optimize each SU(4) gate based on the cost function, e.g. Eq. 10, and then compile each SU(4) gate into native gates of the Quantinuum H1 processor before executing. The energy of the spin system is calculated from sampling correlation functions in the real experiment. See also Supplementary Note 1 for more experimental details.
@@
Here $E$ is a complex variable we optimize over. When $\mathcal{L}^{\prime}=0$ we are guaranteed that |$\ket{\psi({\bm{\theta}})}$ is a right eigenstate of $H$ and $E$ is an eigenvalue (for the rest of the paper, we focus on the property of right eigenstates, and the left eigenstates can be found by substituting $H\leftarrow H^{\dagger}$). Using this method, the authors devised a numerical optimizer and were able to numerically compute the left and right eigenstates, verifying the biorthogonal relations, as well as evaluating many observables [45].
@@
Although the variance minimization scheme presented in Ref. 45 includes all the fundamental components of a variational quantum algorithm, the simulated system sizes were limited to only 7 qubits due to two main obstacles: First, they used a direct, full-state preparation method that is demanding on both classical and quantum memory. Capturing long-range correlations would then require a large circuit volume, making it impractical for larger systems. Second, the termination condition for variance minimization only ensures that some eigenstate is found, but not necessarily the one of interest. Given that the number of eigenvalues increases exponentially with system size, a random guess for $E$ is likely to result in an undesired eigenstate.
@@
Figure 3: (a) The real and (b) imaginary parts of variational right ground state preparation results of a $n=19$ non-Hermitian Ising chain are compared with that from a $\chi=100$ NH-DMRG. The exceptional point can be read from the imaginary part of $E$. We experimentally execute the $J=-1,\kappa\in\{0.2,0.5\}$ circuits for 2,000 measurement shots and report the energy outcomes, which are both within 2% error of the DMRG values. (c) The absolute value of relative energy difference. Notice that the peak on each curve corresponds to the EP read from (b). (d) and (e) repeat the study of (a) and (b), except now the system size is set to 20 so that the system is EP-free. (f) Comparing $ZZ$ correlation functions measured in Quantinuum’s H1 processor (2,000 shots) to DMRG results. All results are prepared with a $q=2$, $\tau=2$ ladder architecture as defined in Fig. 2, and error bars represent one statistical standard deviation. More experimental details on the implementation can be found in the Supplementary Note 1.
@@
To solve the first obstacle, we make use of quantum circuit tensor network state (qTNS) techniques [28], to sequentially simulate many-body quantum states, as shown in Fig. 2. Rather than a direct one-to-one encoding between a spin and a qubit, sequential simulation introduces $q$ ancilla qubits, or bond qubits, with local circuit depth $\tau$ to faithfully capture the near-area law entanglement structure of a physical state, generating a sequence of quantum operations that allows one to sample properties of the many-body state along a spatial direction without storing the full state in quantum memory; we defer a more detailed explanation to Supplementary Note 1.
@@
For the second obstacle, we designed a ‘warm start’ method, as widely used in many other variational algorithms [29]. Namely, we first turn off the non-Hermitian field and solve for the Hamiltonian’s ground state with a VQE algorithm and gradually turn on the imaginary field, feeding the previous optimization result as an initialization. Concretely, we consider the $1d$-transversal field Ising model with an imaginary longitudinal field [46]:
@@
where $J$ and $\kappa$ are real numbers. To find the ground state at a certain $\kappa_{\rm targ}$, we execute the procedure in Algo. 2: Box 2 – VQE algorithm for dissipative Ising model with variance minimizaiton 1:Set the target Hamiltonian to be $H=H_{\rm Ising}(J,\kappa=0)$ 2:Initialize |$\ket{\psi({\bm{\theta}})}$ where the circuit with a variational circuit generated by quantum matrix product state (qMPS) [31] 3:Find the Hamiltonian’s ground state by minimizing $\mathcal{L}=\innerproduct{\psi({\bm{\theta}})|H}{\psi({\bm{\theta}})}$ and record the optimized circuit parameter $\bm{\theta}^{*}$ and energy $E^{*}$; 4:for $\kappa=\Delta\kappa,2\Delta\kappa,\ldots\kappa_{\rm targ}$ do 5:     Initialize the circuit at $\bm{\theta}_{0}\leftarrow\bm{\theta}^{*}$ and $E_{0}\leftarrow E^{*}$ 6:     Minimize the target function $\bra{\psi({\bm{\theta}})}(H-E)(H^{\dagger}-E^{*})\ket{\psi({\bm{\theta}})}$ over $\bm{\theta}$ and $E$. 7:     Record optimized parameter $\bm{\theta}^{*}(\kappa)$ and energy $E^{*}(\kappa)$ 8:end for The dissipative Ising model exerts $\mathcal{PT}$-symmetry, but it goes through spontaneous symmetry breaking at $\kappa_{c}$, the so-called exceptional point: at $\kappa<\kappa_{c}$, the ground state energy remains real despite the non-Hermicity; at $\kappa>\kappa_{c}$, the ground state energy merges with the first excited state to form a complex conjugate pair in their energy values. The exceptional point is very sensitive to $J$ and the parity of the length of the chain.
@@
Additional numerical results for a non-Hermitian XXZ chain with size $n=64$ can be found in the Supplementary Note 1. In Hermitian physics, it is known that quantum computers offer a “data compression” advantage in representing physical states [32, 31, 40], as a quantum computer can store a MPS with bond dimension $\chi$ with merely $\sim\log\chi$ bond qubits. Our result extends this quantum advantage to non-Hermitian quantum material simulations.
@@ III Discussion
We have combined two quantum state compression techniques, GMPS and VQC, to experimentally study the quenched dynamics of an extended strongly-correlated non-Hermitian system. Using a trapped-ion implementation on the Quantinuum H1 quantum processor, we have observed a supersonic mode in the connected density-density correlation function of a fermionic chain following a non-Hermitian quantum quench—a phenomenon conventionally forbidden by the Lieb-Robinson bound in Hermitian systems. The system sizes, time scales, and low sample complexity achieved in our experiments are enabled by the convergence of the advantages offered by universal quantum computers and the variational formulation of the dynamics employed in our experiments, allowing us to bypass costly postselection-based implementations with an affordable approach based on unitary dynamics. Similarly, we exploited the efficient compression offered by qMPS to experimentally prepare eigenstates of a non-hermitian spin chain where we experimentally measured correlation functions and energies across an exceptional point on a dissipative spin chain up to length $n=20$ using only 3 qubits. These results show that universal quantum computers have great potential and offer an advantage as compared to traditional non-universal experimental platforms in the study of non-Hermitian quantum matters.
@@
Our work also raises a number of intriguing questions regarding the opportunities and challenges of using digital quantum computers for simulating non-Hermitian systems. In particular, in Thm. II.1 we established the hardness of the simulation of quench dynamics of certain initial states under single qubit non-Hermitian dynamics. A key open question arising from our experiments and theoretical findings is to determine the properties of the initial state and the non-Hermitian Hamiltonian that enable efficient unitary simulation. Here, a natural direction is to explore the use of VQC to alleviate the postselection requirements which are often encountered in systems such as those exhibiting measurement-induced phase transitions [48]. Similarly, these and other related questions can be reframed in the context of quantum algorithms for solving dissipative differential equations. Specifically, under what conditions can quantum computers provide a significant advantage in solving these equations? This problem has been studied from an algorithmic perspective [49] and can potentially offer new insights into the question of simulability of non-hermitian systems more broadly.
@@
A natural extension of our work is to explore dynamics and eigenstate properties beyond one-dimensional systems where interesting non-Hermitian phenomena emerge [50]. One possibility is, that instead of doing a sudden quench, one could use our protocol to simulate non-adiabatic passage across an EP and investigate the proposed unconventional Kibble-Zurek mechanism [51, 52]. Another interesting direction would be exploring the asymptotic quantum memory costs to approximate the ground states in different non-Hermitian phases of matter. In Hermitian systems, one could relate the hardness of compression to entanglement entropy laws [53]. It is worth pursuing what metric one should use in the non-Hermitian case, as there does not exist a good entanglement measure due to the lack of an orthonormal basis [6]. As we push the boundaries of quantum computation, the techniques and insights gained from our study not only illuminate new pathways for exploring non-Hermitian physics but also pave the way for future advancements in quantum simulation aided by the convergence of innovative techniques such as compression and variational compilation with cutting-edge quantum hardware.
IV Methods
IV.1 Non-Hermitian physics: a motivating example
We argued in the main text that a physical way to motivate non-Hermitian physics is from open quantum system dynamics. Usually, the Linbladian equation is used to describe the Markovian evolution of a system interacting with a thermal bath [54]:
$=-i(H_{\rm eff}\rho-\rho H_{\rm eff}^{\dagger})+\sum_{i}\gamma_{i}L_{i}\rho L^{\dagger}_{i}.$ (13)
The last term in Eq. 40, $\sum_{i}\gamma_{i}L_{i}\rho L^{\dagger}_{i}$, describes quantum jump processes that can be associated with a measurable physical quantity, such as a spontaneous emission of a photon. The stochastic time evolution can be classified into different quantum trajectories depending on the number of quantum jumps. Now, if we postselect on the absence of quantum jumps, we end up with a Schrodinger-like time evolution with a non-Hermitian Hamiltonian:
$\frac{d\rho}{dt}$ $=-i(H_{\rm eff}\rho-\rho H_{\rm eff}^{\dagger})$ (14)
IV.2 Review of sequential circuits
Definition IV.1 ($\tau$-sequential quantum circuit).
  1. 1.
    Strictly contains all constant depth circuits;
  2. 2.
    Is strictly contained in the set of linear depth circuits.
Matrix product states
The discussion of sequential circuits originated from a celebrated quantum state compression method: MPS. Any pure quantum state $|\Psi\rangle$ can be expressed as a matrix-product state by sequentially performing Schmidt decompositions between local sites, turning wave-function amplitudes into a 1D tensor train (in physical states considered in this work, we focus on open boundary 1D chains with sites $x\in\{1,2,\dots n\}$, although the method can be generalized to infinite systems and larger on-site dimensions):
$|\Psi\rangle=\sum_{\{s_{x}\}_{x=1}^{n}}\ell^{T}A^{s_{1}}A^{s_{2}}\dots|s_{1}s_{2}\dots\rangle.$ (15)
Sequential quantum circuits generated by MPS (qMPS)
Box 3 – Generating sequential quantum circuits from MPS 1:Prepare the bond qubits in a state corresponding to the left boundary vector $\ell$. This can be done with up to $\log\chi$ ancilla qubits 2:for $x=1\dots n$ do 3:     Perform a synthesized quantum circuit $U_{A}$ at site $x$, entangling the physical and bond qubits. 4:     Measure the physical qubit in the eigenbasis of $O_{x}$ and weight the measurement outcome by the corresponding eigenvalue of that observable. 5:     Reset the physical qubit for site $x$ in a fixed reference state, $|0\rangle$. 6:end for 7:Discard the bond qubits.
Gaussian MPS
$H=\sum_{i,j=1}^{n}c^{\dagger}_{i}h_{ij}c^{\vphantom{\dagger}}_{j}$ (16)
The compression scheme presented by Fishman and White [39] exploits this unitary invariance by progressively disentangling local degrees of freedom in blocks of $B$ adjacent sites. Moreover, the ground states of Gaussian fermionic systems have near-area-law entanglement. Therefore, choosing block size $B$ large enough, most of the block eigenvalues must be exponentially close to 0 or 1. This enables a sequence of operations that compresses the correlation matrix: Box 4 – GMPS compression 1:for $x=1\dots n$ do 2:     Start with $G_{xx}$, examine the next $B\times B$ sub-matrix of $G$ to the bottom right. 3:     Identify the eigenvector with an eigenvalue closest to 0 or 1. 4:     Apply a series of $2\times 2$ single-particle unitary rotations $\prod_{\alpha=1}^{B-1}u_{x,\alpha}^{\dagger}$ on $G$ to move this eigenvector to the first site of the block. Denote the resultant correlation matrix as $G^{\prime}$. 5:     Set $G\leftarrow G^{\prime}$ 6:end for
GMPS as a sequential quantum circuit
$U_{x,\alpha}=\exp[\sum_{ij}c^{\dagger}_{i}(\log u_{x,\alpha}^{\dagger})_{ij}c^{\vphantom{\dagger}}_{j}]=\exp[\sum_{ij}\sigma^{+}_{i}(\log u_{x,\alpha}^{\dagger})_{ij}\sigma^{-}_{j}].$ (17)
IV.3 Analytical proof of Thm. I.1 and implication in complexity theory
  1. 1.
  2. 2.
The task can be formalized into a decision problem:
Problem IV.2 (Distinguishes a Haar random state from a maximally mixed state).
Lemma IV.3 (Single qubit dissipation distinguishes a Haar random state from a maximally mixed state).
Proof.
$p_{s}:=\left|\innerproduct{s}{\psi_{\rm Haar}}\right|^{2},\ q_{s}:=\left|\innerproduct{s}{\psi_{\rm Haar}(t)}\right|^{2}$ (19)
[…]
$\left(\frac{e^{2t}}{e^{-2t}+e^{2t}}\right)^{n}=c$ (20)
and solving for $t$, we get
$t$ $=-\frac{1}{4}\log(-c^{-1/n}(-1+c^{1/n}))$ (21)
$=-\frac{1}{4}\log(c^{-1/n}-1)$ (22)
$=-\frac{1}{4}\left[\log(-\frac{1}{n}\log(c))+O(n^{-1}))\right]$ (23)
$=\Theta(\log(n))$ (24)
$q_{1^{n}}$ $=\frac{p_{1^{n}}e^{2n}}{\sum_{s}p_{s}e^{4w_{s}-2n}}$ (25)
$=\frac{p_{1^{n}}e^{2n}}{\sum_{i=0}^{n}\sum_{w_{s}=i}p_{s}e^{4i-2n}}$ (26)
$:=\frac{p_{1^{n}}}{p_{1^{n}}+2^{-n}\eta}$ (28)
In the second from last step, we use the fact that each entry of a Haar state vector consists of two standard normal variables and the weights on each string should follow the Porter-Thomas distribution: ${\rm PT}(p_{s})=2^{n}e^{-2^{n}p_{s}}$, and a sum over $\operatorname{poly}(n)$ terms would converge to its mean value. We may also ignore the $p_{0^{n}}e^{-2n}$ in the denominator as it is exponentially small in $n$.
Under the assumption of Eq. 44, we have
$\frac{1}{1+\eta}=c$ (29)
W.l.o.g., setting $c=1/2$ gives $\eta=1$; therefore the Eq. 52 returns
$q_{1^{n}}=\frac{2^{n}p_{1^{n}}}{2^{n}p_{1^{n}}+1},$ (30)
$P[q_{1^{n}}\geq 0.6]$ $=P[q_{1^{n}}\geq 0.6]$ (31)
$=P\left[\frac{2^{n}p_{1^{n}}}{2^{n}p_{1^{n}}+1}\geq 0.6\right]$ (32)
$=P\left[p_{1^{n}}\geq 1.5\times 2^{-n}\right]$ (33)
[…]
$\approx 0.223$ (35)
As evident from Fig. 5, with constant probability, a measurement in the computational basis now allows efficient distinguishing between the Haar random state and the maximally mixed state after the time evolution. Notice that, if we choose $t=\operatorname{poly}(n)$, the two states are once again hard to be distinguished because they both purify to a product state. This probability can be boosted to $1-e^{-k}$ by randomly applying a layer of $X$ gates before the time evolution (namely, randomly selecting a string $s$ to amplify and probe its amplitude) and repeating $O(k)$ times. ∎
[…]
Then the strong state complexity is defined as:
Definition IV.4 (Strong state complexity).
Proposition IV.5 (Strong state complexity of Haar random states).
$\Pr[\mathcal{C}_{\delta}(\ket{\psi_{\rm Haar}})\leq r]\leq 4.0144d(n+1)^{r}|\mathcal{G}|^{r}\exp{-\frac{d(1-\delta^{2})}{9\pi^{3}}}$ ) (38)
Lemma IV.6 (Hardness of distinguishing a Haar random state with a $\mathsf{BQP}$ machine).
Combining Lem. A.3 and Lem. A.6 completes the proof of Thm. I.1. Further, as we explain in Supplementary Note 1, each single-qubit dissipation channel can be implemented with a two-qubit gate and post-selection on the single ancillary qubit. It turns out that the result we have proved can be summarized in a computational complexity language: With overwhelming probability, Prob. A.2 can be solved by a quantum circuit at merely constant depth when postselection is allowed (denoted as $\mathsf{PostQNC_{0}}$), but also with overwhelming probability, it cannot be solved by a polynomial-size quantum circuit. Namely,
Corollary IV.7 (Oracle separation).
There exists a quantum oracle $\mathcal{O}$ (as defined in Prob. A.2) such that $\mathsf{PostQNC_{0}}^{\mathcal{O}}\not\subseteq\mathsf{BQP}^{\mathcal{O}}$; relative to the same oracle, $\mathsf{BQP}^{\mathcal{O}}\subset\mathsf{PostBQP}^{\mathcal{O}}$.
V Data Availability
The generated in this study have been deposited in [72].
VI Code Availability
The simulation code used in this work can be found at [72].
VII Acknowledgement